OpenStudy (anonymous):

Solve by using the quadratic formula x^2 + 6x = -3 -3±√-3?

OpenStudy (campbell_st):

rewrite the equation x^2 + 6x + 3 = 0 then the gneral quadratic formula is \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = 1, b = 6 and c = 3 just substitute and evalaute

OpenStudy (anonymous):

x^2 + 6x = -3 x^2 + 6x + 3 = 0

OpenStudy (anonymous):

-3±√6?

OpenStudy (anonymous):

@campbell_st am i right?

OpenStudy (campbell_st):

looks correct to me... I just did it as mental arithmetic...

OpenStudy (anonymous):

:) thanks